OpenStudy (anonymous):

Find f(x) and g(x) so that the function can be described as y = f(g(x)). y =(4/x^2) + 9

OpenStudy (michele_laino):

I think that a possible choice, can be this: \[\begin{gathered} g\left( x \right) = {x^2} \hfill \\ f\left( x \right) = \frac{4}{x} + 9 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

we have: \[\Large f\left( {g\left( x \right)} \right) = \frac{4}{{g\left( x \right)}} + 9 = ...?\]

OpenStudy (anonymous):

my choices are A.f(x)= x+9 g(x)= 4/x^2 B.f(x)= x g(x)= (4/x)+9 C.f(x)= 1/x g(x)= 4/x)+9 D.f(x)= 4/x^2 g(x)= 9

OpenStudy (michele_laino):

I consider case A: \[\Large f\left( {g\left( x \right)} \right) = g\left( x \right) + 9 = ...\] please substitute g(x9= 4/x^2, what do you get?

OpenStudy (michele_laino):

oops..g(x)=4/x^2, what do you get?

OpenStudy (anonymous):

would it be f(x)= x+9

OpenStudy (michele_laino):

yes! it is, so we can write: \[\Large f\left( {g\left( x \right)} \right) = g\left( x \right) + 9\]

OpenStudy (anonymous):

ohh okay!

OpenStudy (anonymous):

and how would i get a y value?

OpenStudy (michele_laino):

you have to replace g(x) with its definition 4/x^2

OpenStudy (anonymous):

f(g(x))=g(x)+9 = f(g(x))=(4/x^2)+9

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

thank you so much!

OpenStudy (michele_laino):

:)